# Secure and Efficient Migration of Large Enclaves in a Data Center

### v4

@@ 4. Design
In this section, we discuss the design of ConstMig. As shown in Figure 4a, ConstMig handles the migration of the secure portion of the application. CRIU handles the migration of the untrusted portion. The primary design goals of ConstMig are as follows:
@@
Invariant 4: The migration key ($M_{k}$) is a function of the PUF of the source machine, boot time, current time, and the output of a TRNG. A replay attack is thus not possible (§ 6).
@@ 4.1. Threat Model
We make the same security assumptions as Intel SGX and prior work in this area (see Table 1). The operating system and other software entities executing outside the enclave are not trusted. They never see the enclave’s data in plaintext or have access to the keys used during migration. For SGX, the OS handles the page table of a secure application. A malicious OS can manipulate the PRESENT bit in the page table entries (PTEs) to capture the complete page-access pattern of an application (Xu et al., 2015). Such attacks are not in the purview of SGX and are also out of the scope of this work (see §6). We abort the migration if we detect a security violation during it. If $M_{k}$ was already delivered, both enclaves are lost; otherwise, only the restored enclave is affected.
@@ 4.2. High-Level Overview
There are two main components in ConstMig: the migration manager that orchestrates the entire migration process and a trusted SGX library (ConstMigLib), which is linked to the secure application. Its routines have access to the secure memory space of the application running in the enclave.
Start migration: As shown in Figure 4b, at the beginning of the migration, ❶ the migration manager pauses the application and sends the save signal to ConstMigLib. The migration manager then resumes a single thread for ConstMigLib to do its work ($save\,\,thread$). ❷ ConstMigLib generates a master key $M_{k}$ to be used during migration. It then saves the data segment, BSS segment and OptMgr metadata (information about the allocated regions) in an untrusted metadata buffer (MBuff). ❸ A thread starts to save heap data in a separate untrusted backup buffer (BBuff). All these operations are done after performing encryptions and adding integrity checks (§4.5 and §4.6).
@@
Start Restore: ❻ ConstMigLib on the destination machine, decrypts MBuff using $M_{k}$ and reads the encrypted BSS segment, data segment, and memory manager’s metadata. ❻ The migration manager on the destination machine sends the restore signal to ConstMigLib. ❼ A background thread starts reading data from BBuff and starts restoring the enclave’s heap ($restore\,\,thread$). Since BBuff is stored in the untrusted region, it is protected by userfaultfd (Doc, 2022), and the restore thread is guaranteed to read valid pages.
Cleanup: Once the restoration is done, ConstMigLib zeros all the allocated pages on the source machine and kills the application – preventing forking attacls (§6). At any given time, an application runs on only one system.
@@ 4.3.1. Migration Manager
We use the widely popular Checkpoint/Restore In Userspace or CRIU (CRIU Project, 2025; Venkatesh et al., 2019; Stoyanov and Kollingbaum, 2018; Nakashima and Kourai, 2021) package as our migration manager. CRIU is not a part of the trusted codebase, and never sees the migrating enclave data in plaintext. CRIU supports both the pre-copy and post-copy live migration methods. We modify CRIU to handle the SGX enclave migration as follows. A save signal to an enclave triggers a state save process at the beginning of the migration. Conversely, a restore signal triggers restoration when resumed on the destination machine. The signals are captured and processed by ConstMigLib.
@@
Encrypted buffers: We use two data buffers: a metadata buffer (MBuff) to store the BSS and data segments and ConstMig metadata, and a backup buffer (BBuff) to save its heap contents allocated in the untrusted memory region (outside the SGX enclave). MBuff, due to its relatively small size, is transferred synchronously to the destination at the beginning of the migration. BBuff, due to its large size (heap contents), is sent asynchronously. We leverage the userfaultfd mechanism via CRIU for BBuff. Since BBuff is allocated in the untrusted region, a read operation on this buffer on the destination machine will always read the data as it is on the source machine – either the data has already been transferred, or it is pulled synchronously via a network page fault (Hines et al., 2009).
@@
A consistency checker: CRIU handles the restoration of BBuff on the destination machine. However, CRIU is not SGX migration aware. We need to ensure that ❶ BBuff is transferred only after the enclave state has been successfully saved on it, and ❷ the restored enclave only reads restored data. Our consistency checker ensures this by maintaining three bit-vectors: internal save_vec, save_vec and restore_vec of size equal to the number of 4 KB pages in BBuff. Internal save_vec is allocated in the trusted region that indicates when a 4 KB page is ready to be transferred to the destination machine. A copy, save_vec, is also maintained in the untrusted region as it is required by CRIU. restore_vec is stored inside the enclave and indicates whether a page has been restored in the enclave’s memory or not. All entries in save_vec and restore_vec are initially set to zero.
Access tracking: As mentioned earlier (§3.3), in an asynchronous migration, the migration scheme must capture page accesses in a restored application to serve the correct page from the source machine. However, such support is missing for SGX (§3.3.2). In Section 5, we introduce two novel solutions to handle page faults of the restored enclave.
@@ 4.4. ConstMig library (ConstMigLib)
We statically link each SGX application with a trusted library (ConstMigLib). The main components of the library are as follows:
1. Memory Manager: We implement a memory manager (OptMgr) by overriding the default implementations of malloc() and free() to handle all the allocation and free requests. For every allocation request, we store the address of the allocated region, its size, and a unique ID in an array-like structure (MemArr). A free call updates the MemArr by deleting any freed region. During migration, MemArr is encrypted and stored in MBuff and sent synchronously along with the data and bss segments of the enclave to the destination machine.
2. Save thread: A save thread (save_th) handles the save command initiated by CRIU and starts saving the heap’s contents in BBuff. The thread operates at a granularity of 4 KB pages. Upon saving a page, it sets the corresponding bit in save_vec to indicate that the particular page in BBuff is ready to be transferred to the destination machine.
3. Restore thread: A restore thread (restore_th) handles the restore command initiated by CRIU and starts restoring the heap contents from BBuff. It also operates at the granularity of 4KB pages. Upon restoring a particular page, it sets the corresponding bit in the restore_vec to indicate that it has been restored.
4.5. Saving Data
Figure 5. Asynchronous data saving.
ConstMigLib employs a post-copy-based technique to transfer heap data between the source and destination machines. As shown in Figure 5, for each page $P$ at virtual address ${VA_{p}}$, save_th performs the following actions: concatenates the page contents ($PC(P)$) with the address ($PC(P)||VA_{p}$) and calculates its hash ($H$). It then concatenates $PC(P)$ with the hash and encrypts both using $M_{k}$. The encrypted packet is stored in BBuff, and the corresponding bits in save_vec and internal save_vec are updated.
4.6. Restoring Data
At the destination node, ConstMigLib restores data using restore_th (see Figure 6). This thread retrieves encrypted data from the BBuff, decrypts it within SGX using $M_{k}$ and extracts the page contents ($PC(P)$) and hash ($H$). It then calculates the expected hash ($H^{\prime}$) and validates that $H==H^{\prime}$. If validation is successful, it stores the data on the corresponding page; otherwise, a security violation is reported, and migration is aborted (see Figure 6)). After restoration, restore_th sets the corresponding bit in restore_vec, marking the page that has been restored. If the bit in restore_vec was already set, a security violation is reported, and migration is aborted (see Invariant 4).
@@ 4.6. Restoring Data
4.6.1. Network Page Fault
If the requested page is not available in BBuff, CRIU issues a network page fault request to the source machine to retrieve the page and restore_th enters a blocked state. Upon receiving the request, the source machine’s CRIU signals save_th to save the page into BBuff. save_th first checks the internal save_vec data structure for the availability of the page in BBuff. A security violation occurs if the page is already saved. save_th then encrypts the page and stores it in BBuff4.5). The encrypted page is then sent to the destination machine by CRIU. Upon receipt, ConstMigLib validates the encrypted contents (§4.6).
@@ 5. Fine-Grained Access Tracking
A key challenge in migration is that the restored enclave and the SGX hardware are not migration-aware. The restored SGX enclave on the destination machine can access any memory region within its address space. If it attempts to access a memory location that is yet to be restored by restore_th, the enclave will fault and will be allocated a fresh zeroed data page by the fault handler instead of the page from the source machine (present in the encrypted form as a page in BBuff). We call this issue as use-before-restore. As discussed in Section 3, the userfaultfd mechanism that ensures that the data in BBuff is consistent does not work for a memory region inside an enclave. Page faults in the SGX secure region cannot be serviced using userfaultfd as it does not support device-backed memory pages (SGX pages are backed in the device “/dev/isgx” using the default SGX driver) in the Linux kernel (Corbet, 2022, 2023; Kernel, [n. d.], 2023).
@@
To address this issue, we propose two novel schemes: the ConstMig fault tracker and the ConstMig access tracker. They ensure that a restored enclave reads valid data instead of a fresh all-zero page.
(a) Fault Tracker
(b) b
Figure 7. Two different approaches for fine-grained access tracking: (a) using page faults or (b) using automatically added access checks. NF: Network faults.
5.1. ConstMig Fault Tracker
The first approach extends the SGX page fault handler to serve restored data instead of a fresh page on the destination machine. It leverages the fact that the operating system manages the page table of an enclave. A page fault is triggered if the application attempts to access a page that is yet to be allocated. The SGX fault handler services the fault and returns a zeroed page to the enclave. We modify the SGX fault handler to return a “restored” page instead of a zeroed page. Note that the SGX fault handler executes as a kernel module and can only read the encrypted data stored in BBuff. The fault handler must use the services from the enclave itself to decrypt, validate and restore the data at the correct location.
@@ 5.1. ConstMig Fault Tracker
Fault capture and restoration: As shown in Figure 7a, we modify the SGX driver ❶ to send a signal to the application on a fault. ❷ Upon receiving the signal, the application makes an ECALL into the ConstMigLib enclave with the faulting address. ConstMigLib checks if the page is already restored using restore_vec. If the page is already restored, it simply returns. ConstMigLib also checks if restore_th is currently restoring this page and if the fault is due to that. ❸ If the fault is not due to restore_th, ConstMigLib reads the encrypted page from BBuff (this, in turn, may cause a network page fault that is then transparently handled by CRIU), it then decrypts the page, validates its integrity, and ❹ writes the decrypted data to the faulting address. Now, the original faulting thread will see the restored data instead of all zeros. Note that there will be exactly one signal for every page in the enclave either due to restore_th or a page fault in the enclave itself (Invariant 3).
5.1.1. Implementation:
We use the protection bits in the page table entries (Gorman, 2012) to raise a fault whenever the restored enclave tries to access a page. We set the PROT_NONE bit for all the committed pages using the mprotect call (mpr, [n. d.]). Specifically, we modify the sig_handler function in the trusted part of the SGX SDK (called PSW (Intel, 2019a)) to handle protection faults and send a signal to the application.
The signal is sent after the protection bit is cleared in the page table entry. We implement a dedicated trusted thread in SGX SDK called constmig_fault_thread which makes the ECALL from the application to ConstMigLib upon receiving a page fault signal. The ECALL is required to decrypt, check the integrity, and restore the data. A separate thread is needed because the thread that caused the page fault earlier is still waiting for the fault to be resolved, and using it will crash the enclave.
@@ 5.1.1. Implementation:
5.2. ConstMig Access Tracker
The second approach is based on the idea of adding a “check” before every heap pointer access in the restored enclave (see Figure LABEL:fig:access_tracker). The check uses the information stored in MemArr to ensure that the pages about to be accessed by the enclave have been restored from BBuff. A sample code that shows the added check to restore a page using a function called __process_access() is shown in Listing 1. Note that the check is executing within the enclave and does not require an ECALL.
The function __process_access() takes two arguments: the pointer address from which the access will start and the access size. It does a linear scan of MemArr to find the memory region in which the current access request belongs. A linear scan is suitable when there are only a few allocated regions. However, a radix tree-like structure can facilitate efficient lookups for a large number of memory regions. Upon finding the correct memory region, it finds the pages in BBuff that must be restored to service the request using the location of the memory region in the BBuff, its address, and the parameters of the memory access request.
It then checks restore_vec to see if all the required pages have already been restored. If all of them have been restored, it returns; otherwise, it reads the encrypted pages from BBuff and restores the missing pages. Note that reading the encrypted pages in BBuff might raise network faults, which are transparently handled by CRIU. Once the restoration of all the pages is done (indicated by a global flag), the memory check is replaced with a return 0; statement.
@@ 5.2. ConstMig Access Tracker
Implementation: We implement a Clang-based source-to-source transformer (Bendersky, [n. d.]; cla, [n. d.]) to add additional memory checks. This is a common approach to ensure the validity of a pointer before dereferencing it and has been used extensively (Nagarakatte et al., 2009; Richardson, 2020; Nagarakatte et al., 2015; Saileshwar et al., 2022; Richardson, 2020). Our transformer extends the RecursiveASTVisitor class in Clang to traverse the abstract syntax tree (AST) (ast, [n. d.]). An AST is a tree representation of the constructs (variables, functions, etc.) in the source code. We use the AST tree to get information about heap variable usage. The RecursiveASTVisitor class in Clang offers standard APIs to process different kinds of nodes in the AST. For example, the heap variable pos in Listing 1 shows up as a DeclRefExpr node in the AST. DeclRefExpr contains the name of the variable and also its type. Our transformer uses this information to add the __process_access() check before accessing the variable.
5.3. Discussion
The choice between the two trackers depends on the extent of changes a cloud vendor is able to implement on its system. Fault Tracker requires no enclave source code modification (apart from adding a signal handler in the untrusted part of the application); however, it is expensive because it involves sending a signal and making an ECALL. On the other hand, Access Tracker requires enclave source code modification. Its migration is less overhead-prone because it does not require any signals or ECALLs (as the checks are done inside the enclave). The cost to restore a page (read from BBuff and decryption) is the same for both approaches. After migration, Access Tracker has a small runtime overhead, which Fault Tracker does not have (§7.4.1).

### v5

@@ 3. Design
In this section, we discuss the design of ConstMig. As shown in Figure LABEL:fig:scoping, ConstMig handles the migration of the secure portion of the application. CRIU handles the migration of the untrusted portion. The primary design goals of ConstMig are as follows:
@@
Invariant 4: Migration components executing outside SGX never see the plain text data.
Invariant 5: The migration key ($M_{k}$) is a function of the PUF of the source machine, boot time, current time, and the output of a TRNG. A replay attack is thus not possible (§ 5).
3.1. Scope
As discussed earlier, modern TEEs support hardware-assisted live migration, with the notable exception of Intel SGX. Given Intel’s plans to continue SGX support on future server-class processors, enabling live migration for large SGX enclaves remains an important and timely challenge. Our focus is primarily on heap memory, which constitutes the majority of an application’s memory footprint (Aga and Narayanasamy, 2019). In contrast, other segments such as data and BSS typically range from a few KBs to MBs and are always transferred synchronously as part of the minimal state during migration.
Furthermore, we do not discuss migration of sealed data. SGX provides a way to seal data using a hardware-bound key. Alder et al.(Alder et al., 2018) proposed a method to migrate sealed data by encrypting all the sealed data with a software-generated key on the source machine and resealing it on the destination machine after transfer and decryption. This solution can be easily incorporated into ConstMig.
@@ 3.2. Threat Model
We adopt the same security assumptions as Intel SGX as prior work in this domain. The operating system and other software components outside the enclave are not trusted. They do not have access to enclave plaintext data or the cryptographic keys used during migration.
In SGX, the OS manages the page tables of secure applications. A malicious OS can manipulate the PRESENT bit in page table entries (PTEs) to infer the application’s page access patterns (Xu et al., 2015). Such side-channel attacks fall outside the threat model of SGX and are beyond the scope of this work (see §5). Migration is aborted immediately upon detection of any security violation.
@@ 3.3. High-Level Overview
There are two main components in ConstMig: the migration manager (CRIU) that orchestrates the entire migration process and a trusted SGX library (ConstMigLib), which is linked to the secure application (see Figure LABEL:fig:scoping). Its routines have access to the secure memory space of the application running in the enclave.
Start migration: As shown in Figure LABEL:fig:optmig_high_level, at the beginning of the migration, ❶ the migration manager pauses the application and sends the save signal to ConstMigLib. The migration manager then resumes a single thread for ConstMigLib to do its work (savethreadsave\,\,thread). ❷ ConstMigLib generates a master key $M_{k}$ to be used during migration. It then saves the data segment, BSS segment and OptMgr metadata (information about the allocated regions) in an untrusted metadata buffer (MBuff). ❸ A thread starts to save heap data in a separate untrusted backup buffer (BBuff). All these operations are done after performing encryptions and adding integrity checks (§3.5.2 and §3.5.3).
@@
Start Restore: ❻ ConstMigLib on the destination machine, decrypts MBuff using $M_{k}$ and reads the encrypted BSS segment, data segment, and memory manager’s metadata. ❻ The migration manager on the destination machine sends the restore signal to ConstMigLib. ❼ A background thread starts reading data from BBuff and starts restoring the enclave’s heap (restorethreadrestore\,\,thread). Since BBuff is stored in the untrusted region, it is protected by userfaultfd (Doc, 2022), and the restore thread is guaranteed to read valid pages.
Cleanup: Once the restoration is done, ConstMigLib zeros all the allocated pages on the source machine and kills the application – preventing forking attacks (§5). At any given time, an application runs on only one system.
@@ 3.4.1. Migration Manager
We use the widely popular Checkpoint/Restore In Userspace or CRIU (CRIU Project, 2025; Venkatesh et al., 2019; Stoyanov and Kollingbaum, 2018; Nakashima and Kourai, 2021) package as our migration manager. CRIU supports both the pre-copy and post-copy live migration methods. In ConstMig, CRIU is not a part of the trusted code base, and never sees the migrating enclave data in plaintext.
To support SGX enclave migration, we extend CRIU with enclave-aware functionality. A save signal initiates a state save process and marks the beginning of a migration, while a restore signal triggers restoration when resumed on the destination machine. The signals are captured and processed by ConstMigLib.
extbfEncrypted buffers: We utilize two encrypted buffers during migration: a metadata buffer (MBuff) that stores the BSS, data segments, and ConstMig-specific metadata, and a backup buffer (BBuff) to store the encrypted heap contents. Both buffers are allocated in untrusted memory outside the SGX enclave.
@@
Due to its relatively small size, MBuff is transferred synchronously to the destination at the beginning of migration. In contrast, BBuff – containing large encrypted heap data – is transferred asynchronously. We integrate the userfaultfd mechanism via CRIU to support on-demand page retrieval for BBuff. Since it resides in untrusted memory, read operations on the destination machine either access already-transferred data or trigger synchronous page fetches over the network (Hines et al., 2009).
A consistency checker: CRIU handles the restoration of BBuff on the destination machine. However, CRIU is not SGX migration aware. To ensure correctness, we enforce two key constraints:
  1. (1)
    BBuff is transferred only after the enclave state has been successfully saved on it.
  2. (2)
    The restored enclave only reads restored data.
@@
To enforce these constraints, our consistency checker maintains three bit-vectors—internal save_vec, save_vec, and restore_vec of size equal to the number of 4 KB pages in BBuff. The internal save_vec, allocated within the trusted region, tracks when a page is ready for transfer. A copy of this vector, save_vec, is maintained in the untrusted region for CRIU’s use. restore_vec, stored inside the enclave, indicates whether a page has been restored into enclave memory. All entries in save_vec and restore_vec are initialized to zero.
extitFine-grained Access tracking: To ensure that a restored enclave accesses only valid pages, we enforce that a page read is permitted only if the corresponding bit in restore_vec is set. Otherwise, the access is paused until the page is fetched from the source machine. Such support is missing from SGX as of now. We introduce two novel mechanisms to capture enclave memory accesses: access tracker and fault tracker. The access tracker automatically inserts memory access checks into the application’s source code (using an LLVM pass), enabling fine-grained control. In contrast, the fault tracker leverages page table entries to detect access attempts at runtime.
Note that using page table entries to track leaks the memory access pattern. However, as discussed earlier (§3.2), such leakage is unavoidable in SGX, as the OS –an untrusted entity –manages the page tables (§3.2). We provide a detailed comparison of these approaches in Section 4.
@@ 3.5. ConstMig library (ConstMigLib)
(a) Asynchronous data saving.
(b) Asynchronous data restore.
@@
To perform migration, ConstMig requires knowledge of the application’s allocated memory regions, along with mechanisms to save and restore them. We achieve this by statically linking ConstMigLib to the SGX application, enabling support for memory region discovery and migration. ConstMigLib executes inside SGX and has access to all of the application’s address space in plain text form. Its main components are:
3.5.1. Memory Manager
We implement a custom memory manager (OptMgr) by overriding the default malloc() and free() functions to intercept all allocation and deallocation requests. For each allocation, we record the address, size, and a unique ID in a structured array (MemArr). Upon deallocation, the corresponding entry is removed from MemArr. During migration, MemArr is encrypted and stored in MBuff, which is transferred synchronously along with the enclave’s data and BSS segments to the destination machine.
3.5.2. Save thread
A dedicated save thread (save_th) handles the save command initiated by CRIU and begins capturing the heap contents into BBuff. As shown in Figure 6(a), operating at a granularity of 4 KB pages, save_th performs the following steps for each page $P$ at virtual address $VA_{p}$: it concatenates the page contents $PC(P)$ with the address to form $PC(P)||VA_{p}$, and computes a hash $H$ over this concatenation. The page contents and hash are then encrypted using $M_{k}$, and the resulting encrypted packet is stored in BBuff. Once a page is saved, the corresponding bits in both save_vec and internal save_vec are updated to indicate readiness for transfer to the destination machine.
3.5.3. Restore thread
A dedicated restore thread (restore_th) handles the restore command initiated by CRIU and begins restoring heap contents from BBuff. Operating at a granularity of 4 KB pages, the thread retrieves encrypted data from BBuff, decrypts it within SGX using $M_{k}$, and extracts the page contents ($PC(P)$) and hash ($H$). It then computes the expected hash ($H^{\prime}$) and verifies that $H==H^{\prime}$. If validation succeeds, the page is loaded into enclave memory; otherwise, a security violation is reported and migration is aborted (see Figure 6(b)). After successful restoration, restore_th sets the corresponding bit in restore_vec, marking the page as restored. If the bit was already set, indicating a duplicate restoration attempt, a security violation is reported and migration is aborted (see Invariant 3).
Network Faults: During restore, if the requested encrypted page is not available in BBuff (yet to be copied from the source machine), CRIU issues a network page fault to the source machine, and restore_th enters a blocked state. Upon receiving the request, the source machine’s CRIU signals save_th to save the page into BBuff. save_th first checks the internal save_vec to verify whether the page has already been saved. If so, a security violation is reported. Otherwise, the page is encrypted and stored in BBuff3.5.2), and CRIU transfers the encrypted page to the destination. Upon receipt, restore_th validates the encrypted contents before restoring them into enclave memory.
Algorithm 1 Prepare Encrypt Packet
1:procedure PrepareEncryptPacket($P,VA_{P},M_{k}$)
2:  Check(internal save_vec)
3:  $H\leftarrow\text{Hash}(PC(P)\|VA_{P})$ $\triangleright$ Link page content with address
4:  $C\leftarrow\text{Encrypt}(PC(P)\|H,M_{k})$
5:  Update (save_vec, internal save_vec)
6:  return $C$
7:end procedure
Algorithm 2 Validate Encrypt Packet
1:procedure ValidateEncryptPacket($C,VA,M_{k}$)
2:  Check(restore_vec)
3:  $PC(P)\|H\leftarrow\text{Decrypt}(C,M_{k})$
4:  $H^{\prime}\leftarrow\text{Hash}(PC(P)\|VA)$
5:  if $H=H^{\prime}$ then
6:   Update(restore_vec)
7:   return true
8:  else
9:   return false
10:  end if
11:end procedure
@@ 4. Fine-Grained Access Tracking
A key challenge in enclave migration is that both the restored enclave and SGX hardware are not migration-aware. Upon restoration, the enclave can access any addressable memory region; however, if it touches a page not yet restored by restore_th, it triggers a page fault. Instead of retrieving the original encrypted page from BBuff, the fault handler allocates a fresh zeroed page – leading to a use-before-restore issue. As discussed in Section 2.4, this cannot be mitigated using userfaultfd, since SGX pages are device-backed (via /dev/isgx/dev/isgx) and unsupported by userfaultfd in the Linux kernel (Corbet, 2022, 2023; Kernel, ; Kernel, 2023).
@@
To address this issue, we propose two novel schemes: the fault tracker and the ConstMig access tracker. They ensure that a restored enclave reads valid data instead of a fresh all-zero page.
4.1. Fault Tracker
Figure 7. Working of Fault Tracker to ensure a consistent memory access in the restored enclave
The first approach (see Figure 7) modifies the SGX page fault handler to return restored data instead of zeroed pages on the destination machine. Since enclave page tables are managed by the OS, a fault is triggered when the enclave accesses an unallocated page. By default, the SGX fault handler returns a zeroed page; we extend it to fetch the corresponding encrypted page from BBuff (which may cause a network fault that is transparently handled by CRIU). However, the page must be decrypted before the application can use it. As the fault handler operates in kernel space and cannot decrypt enclave data, it invokes enclave services to perform decryption, validation, and placement of the restored page.
@@ 4.1. Fault Tracker
Fault capture and restoration: As illustrated in Figure 7, we extend the SGX driver to signal the application upon a page fault. Upon receiving the signal, the application performs an ECALL into the ConstMigLib enclave with the faulting address. ConstMigLib first checks whether the fault is due to an ongoing restoration by restore_th. If not, it retrieves the encrypted page from BBuff, decrypts and verifies its integrity, and writes the restored data to the faulting address. This ensures that the faulting thread receives the correct data instead of a zeroed page. Importantly, each enclave page triggers exactly one signal – either during restoration or due to a fault – maintaining Invariant 3.
Implementation: To intercept page accesses in the restored enclave, we leverage protection bits in page table entries (Gorman, 2012), setting PROT_NONE on committed pages via mprotect (68). We modify the sig_handler in the trusted SGX SDK (PSW (Intel, 2019a)) to handle protection faults and signal the application after clearing the protection bit. A dedicated trusted thread, constmig_fault_thread, is introduced in the SGX SDK to handle these signals. This thread performs an ECALL into ConstMigLib to decrypt, verify, and restore the faulting page. A separate thread is necessary since the faulting thread remains blocked until the fault is resolved; reusing it would crash the enclave.
4.2. Access Tracker
Figure 8. Working of Access Tracker to ensure a consistent memory access in the restored enclave
The second approach introduces a runtime check before each heap pointer access in the restored enclave (see Figure 8). During an active migration, this check consults restore_vec to verify whether the target pages have been restored from BBuff. Listing 1 shows an example using the __process_access() function, which executes entirely within the enclave and does not require an ECALL. Once migration is complete (all the pages are restored – indicated by a global flag), the memory check is replaced with return 0;.
The function __process_access() takes a pointer address and access size, performs a linear scan of MemArr– the array that contains information about all the allocated regions – to locate the corresponding memory region, and identifies the required pages in BBuff. While linear scans suffice for a small number of regions, a radix tree can optimize lookups at scale. It then checks restore_vec to determine if all required pages are already restored. If not, it reads the encrypted pages from BBuff (may cause network faults, handled transparently by CRIU), decrypts, and restores them.
@@ 4.2. Access Tracker
extbfImplementation: We implement a Clang-based source-to-source transformer (12; 67) to insert memory access checks before heap pointer dereferences. This technique, widely used for pointer safety (Nagarakatte et al., 2009; Richardson, 2020; Nagarakatte et al., 2015; Saileshwar et al., 2022), ensures that accessed pages have been restored. Our transformer extends Clang’s RecursiveASTVisitor to traverse the abstract syntax tree (AST) (1), which represents source code constructs such as variables and functions. For instance, the heap variable pos in Listing 1 appears as a DeclRefExpr node, containing its name and type. Using this information, the transformer inserts a call to __process_access() before each heap pointer access.
4.3. Fault Tracker v.s. Access Tracker
The choice between the two trackers depends on the extent of changes a cloud vendor is able to implement in its system.
Access to Source code Fault Tracker requires no enclave source code modification (apart from adding a signal handler in the untrusted part of the application). On the other hand, Access Tracker requires enclave source code modification to add checks before pointer accesses.
Page restoration Fault tracker restoration is expensive because it involves sending a signal and making an ECALL for each page restoration, whereas Access tracker is less overhead-prone because it does not require any signals or ECALLs (as the checks are done inside the enclave).
Figure 9 shows the time taken by each step in Fault and Access Trackers (resp.) for 4 KB data and 4 MB data, respectively. For 4 KB and 4 MB of data, Access Tracker outperforms Fault Tracker by $35\%$ and $93\%$, respectively. The performance improvement for Access Tracker grows with an increase in the size of the data to be restored. Fault Tracker operates at a 4 KB granularity; hence, restoring 4 MB of data will require $\approx$ 1,024 faults, signals, and ECALLs. Whereas Access Tracker can potentially use the MemArr to restore the complete 4 MB region using a single ECALL. The cost to restore a page (read from BBuff and decryption) is the same for both approaches. Since Access Tracker outperforms Fault Tracker, we use it for the rest of the experiments.
Figure 9. Performance of access tracker and fault tracker approach for consistency checker for 4 KB and 4 MB data. As fault tracker needs to fault on each page to restore it, the overhead of ensuring consistency checker is much higher compared to access tracker. FT: Fault Tracker, AT: Access Tracker
Post Migration After migration, Access Tracker has a small runtime overhead, which Fault Tracker does not have (§6.4).